OpenStudy (anonymous):

s^3t^2/10s^2t^4 ÷ 8s^4t^2/5t^6

zepdrix (zepdrix):

Oh snap, it's varus! :D

zepdrix (zepdrix):

Recall that when we are `dividing by a fraction`, we can rewrite it as `multiplication` by `flipping the bottom fraction.`\[\Large \frac{s^3t^2}{10s^2t^4}\div\frac{8s^4t^2}{5t^6}\quad=\quad \frac{s^3t^2}{10s^2t^4}\cdot\frac{5t^6}{8s^4t^2}\] That part make sense?

OpenStudy (anonymous):

Lol, thanks!